40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3212     Entry time: Wed Jul 14 01:05:27 2010
Author: Sharmila,Katharine 
Type: Summary 
Category: elog 
Subject: Maglev 

Yesterday we hooked up the Quadrant Maglev control to the power supply to test the components in the Input/Output part of the circuit.

The output from the buffer was an unexpected high noise signal which was caused by some circuit components.

Consequently these were replaced/removed after confirming the source of noise.

The following is a story of how it was done.


To test the components of input/output, we measured the output across TP_PD3(Test Point -Photo Diode 3).
We got a high noise signal with a frequency of several kHz.

We tested the values of various electronic components. The resistances R5 and R6 did not measure as mentioned(each had a value of 50 K in the schematic). The value of R6 was 10 K and we replaced R5 with a 10 K resistor. We still got the noise signal at 5.760 kHz with a Pk-Pk voltage of 2.6 V. The resistors in R-LED measured 1.5 K instead of the marked 2.2 K.

P7120278.JPG
We had three suspects in hand:

  • BUF634P : A buffer from the Sallen-Key filter to the LED.
  • C24 : A capacitor which is a part of the Sallen-Key filter.
  • C23 : A capacitor in the feedback circuit of the Sallen-Key filter.


BUF634P : The data sheet for the BUF634P instructed a short across the 1-4 terminals in the presence of capacitive load.  We followed this to overcome the effect(if any) of the extra-long BNC cables which we were using. The oscilloscope still waved 'Hi!' at a few kHz. We removed the buffer and also the feedback resistor R42 from the circuit, what we were testing boiled down to measuring the output of the Sallen-Key filter. The output still contained the funny yet properly periodic signal at a few kHz.      

.P7120284.JPG    


C24: Removing C24 did not do any good.

C23: As a final step C23 was removed. And ... We got a stable DC at 9.86 V(almost stable DC with a low noise at a few MHz). C24 and the buffer were replaced and output seemed fine. The output was a high frequency sine wave which was riding on a DC of 9.96 V.

 

P7120281.JPG


We rechecked if the LED was on and the infrared viewer gave a positive signal.



We went ahead obtaining the transfer function of the feedback control for which we used a spectrum analyzer.

The input for feedback system is a photo current whereas the spectrum analyzer tests the circuit with a voltage impulse.  Hence the voltage input from the spectrum analyzer needs to be converted into current of suitable amplitude(few microamps) for testing the spectrum analyzer.  Similarly the output which is a coil current needs to be changed to a voltage output through a load for feeding into the channel of the spectrum analyzer. We used a suitable resistance box with BNC receiving ends to do this. We obtained a plot for the transfer function which is shown below.

P7140292.JPG


Future plans:

- Check the calculated transfer functions with the plot of the spectrum analyzer

- Model the entire(OSEM, magnet, actuators etc.) system in Simulink and calculate the overall transfer function

- Stable levitation of the 1X1 system

ELOG V3.1.3-